OpenStudy (anonymous):

solve.. x(squared)-6x= -8

OpenStudy (anonymous):

(x-2)(x-4) Do u need working????

OpenStudy (anonymous):

x^2 -6x + 8 = 0 (x-4)(x-2)=0 x = 2 or 4